OpenStudy (anonymous):

PLEASE solve log(4x + 10)= 3.

OpenStudy (weasel123):

@pooja195

OpenStudy (richgall____):

SINCE You know that a positive number won't get you down to 3 you have to find a negative number to multiply by

OpenStudy (anonymous):

-7/4 thank you!

OpenStudy (anonymous):

\[\log(4x+10)=3\iff 4x+10=10^3\]